# LEMON-Mapping: Loop-Enhanced Large-Scale Multi-Session Point Cloud Merging and Optimization for Globally Consistent Mapping

### v1

@@ Abstract
With the rapid development of robotics, multi-robot collaboration has become critical and challenging. One key problem is integrating data from multiple robots to build a globally consistent and accurate map for robust cooperation and precise localization. While traditional multi-robot pose graph optimization (PGO) maintains basic global consistency, it focuses primarily on pose optimization and ignores the geometric structure of the map. Moreover, PGO only uses loop closure as a constraint between two nodes, failing to fully exploit its capability to maintaining local consistency of multi-robot maps. Therefore, PGO-based multi-robot mapping methods often suffer from serious map divergence and blur, especially in regions with overlapping submaps. To address this issue, we propose Lemon-Mapping, a loop-enhanced framework for large-scale multi-session point cloud map fusion and optimization, which reasonably utilizes loop closure and improves the geometric quality of the map. We re-examine the role of loops for multi-robot mapping and introduce three key innovations. First, we develop a robust loop processing mechanism that effectively rejects outliers and a novel loop recall strategy to recover mistakenly removed loops. Second, we introduce a spatial bundle adjustment method for multi-robot maps that significantly reduces the divergence in overlapping regions and eliminates map blur. Third, we design a PGO strategy that leverages the refined constraints of bundle adjustment to extend the local accuracy to the global map. We validate our framework on several public datasets and a self-collected dataset. Experimental results demonstrate that our method outperforms traditional map merging approaches in terms of mapping accuracy and reduction of map divergence. Scalability experiments also demonstrate the strong capability of our framework to handle scenarios involving numerous robots.
@@ I Introduction
Large-scale 3D mapping is a fundamental capability in modern robotics, providing rich geometric information that supports tasks such as drone-based inspection [2, 3, 4], autonomous driving [5], and long-term exploration with ground robots [6]. Furthermore, large-scale 3D maps are also critical for emerging fields such as embodied AI, where agents interact with complex environments based on spatial understanding [7], and end-to-end visuomotor navigation systems [8], which rely on accurate environmental priors to improve generalization ability and robustness. Especially, multi-robot 3D mapping is essential for large-scale and complex tasks, where heterogeneous robot teams provide greater coverage and robustness compared to a single agent. In scenarios such as search and rescue [9], forest monitoring [10], and subterranean exploration [11], accurate point cloud fusion is vital to build consistent and high-quality maps under complex and GPS-denied conditions.
@@
Rapid development of LiDAR-based Simultaneous Localization and Mapping (SLAM) techniques [12, 13, 14, 15, 16, 17] has enabled individual robots to perform real-time localization and point cloud map construction. However, one major challenge remains. Most existing lightweight SLAM systems [14, 15, 16] designed for onboard robot computers are based on direct LiDAR-Inertial-Odometry methods. These methods prioritize pose estimation over map construction to achieve real-time performance. However, the lack of feature extraction and matching in these approaches results in maps with blurred details and poor geometric accuracy. Therefore, the quality of the point cloud maps generated by onboard computers based on lightweight methods requires further improvement.
@@
To address mapping challenges and enhance the geometric accuracy of point cloud maps, various optimization strategies have been proposed. Among them, one category of approaches leverages pose graph optimization (PGO) [18] to handle loop scenarios. Loop closure constraints detected by loop detection methods [19, 20, 21, 22] are added to the pose graph [18] along with odometry constraints to eliminate accumulated drift and improve the global consistency of the map. However, PGO-based methods only use the point cloud map to extract descriptors during loop detection and ignore the structure and geometric information. Therefore, the problem of inaccurate geometric structure and blurred map is not fundamentally solved. (We provide examples in Fig. 8 (b) and (c)). In contrast, another type of map refinement method based on bundle adjustment (BA) has demonstrated impressive capabilities to enhance point cloud maps in recent years [23, 24, 25, 26]. These approaches utilize the structural characteristics of the point cloud map and improve its quality by minimizing the geometric residuals between feature points and feature structure. Nevertheless, the temporal-sequential nature of these methods limits their ability to actively and effectively leverage loop closure information when dealing with scenarios involving multiple revisits to the same area, which may ultimately lead to degraded global consistency (as illustrated in Fig. 7 (b) and (c)). The dependence on the chronological ordering of poses also limits the applicability of BA-based methods to multi-robot systems.
@@
In addition to the above mapping approaches, to achieve larger-scale map reconstruction and multi-robot collaboration, several multi-robot SLAM systems [27, 28] and offline map fusion techniques [29, 30, 31, 32, 33] have shown promising results in recent years. In general, these methods focus on two aspects: removing outliers in loop closure data and achieving map fusion through multi-robot pose graph optimization. However, as discussed in the previous paragraph, PGO-based methods neither directly utilize the geometric information of multi-robot maps nor fully leverage loop closure data. Therefore, these methods only roughly ensure the basic alignment and consistency of the global map, but they fall short in producing high-quality point cloud maps. In particular, they often suffer from blur and misalignment in regions with overlap between robots (as demonstrated in the comparison results in Fig. 8 (b) and (c)).
@@
Inspired by the strengths and limitations of the aforementioned approaches, we conduct an in-depth analysis of previous multi-robot SLAM systems and map merging systems. We argue that the merging of point cloud maps is essentially a map-driven local point cloud registration problem rather than a pose-dominated pose optimization problem. Building upon this perspective, we identify two essential challenges for achieving effective multi-robot point cloud map fusion: (1) Map fusion should be treated as a map-driven registration problem rather than a pose-centric optimization task, and PGO alone can not solve this problem. (2) Loop closure information should be fully exploited in registration beyond time-sequential constraints to ensure both global consistency and local accuracy.
@@
To overcome the above challenge, we propose LEMON-Mapping, a Loop-Enhanced large-scale Multi-session map merging and OptimizatioN framework that achieves globally consistent and geometrically accurate point cloud mapping. Our framework re-examines the ability of loop closure and reasonably enhances its role by two aspects, which effectively improves the consistency of merged map and eliminates the divergence in overlapping area (shown in Fig. 8). First, we innovatively recall loops that are mistakenly removed after obtaining the prior information of multi-robot trajectories, providing more sufficient information and optimization starting points for subsequent spatial bundle adjustment. Second, by leveraging sufficient valid loop closures, our proposed spatial bundle adjustment overcomes the limitation of traditional bundle adjustment methods that operate only on single-robot maps. The spatial bundle adjustment is performed within a local spatial window around the loop, effectively leveraging multiple observations and abundant constraints from different robots. It directly optimizes the poses from different robots equally within the window; thus, the poses of multiple robots matching the same geometric features (planes, lines) are adjusted simultaneously, generating a locally accurate and consistent map. Although spatial bundle adjustment improves local map accuracy at loop closures, it lacks the capability for global map fusion. To address this problem, we propose a reasonable PGO-based way to transfer the local alignment achieved by our spatial BA to the whole map, which effectively combines local BA constraints and odometry constraints to attain a globally consistent and accurate map. We conduct a series of extensive experiments, and the results demonstrate the high mapping accuracy and strong scalability of our multi-session map fusion approach, as shown in Fig. LABEL:fig:headimg and Fig. 13.
@@
  • A scalable multi-session point cloud map merging and optimization framework is designed, which integrates two-stage pose graph optimization with spatial bundle adjustment. The method achieves high-precision 3D mapping reconstruction in large-scale and multi-robot scenarios.
  • A novel spatial bundle adjustment that can be used for multi-robot mapping is introduced, which operates on loop-based spatial windows to fully utilize loop constraints and jointly optimize multi-robot poses equally. This improves the local accuracy and eliminates the serious map divergence.
  • A pose graph optimization scheme with reasonable sparsification of BA constraints is developed, which effectively propagate local accuracy to the global scale, improving both consistency and efficiency.
@@ II Related Works
Figure 1: The framework of our method. (a) The loop processing module. (b) The map Merging module. Our framework takes multi-agent submaps and odometry as input, and generates accurate and globally consistent merging map.
@@ II-A Single Map Maintenance and Optimization
Single map maintenance and optimization have been extensively studied, with existing approaches falling broadly into two categories: pose graph optimization (PGO) and bundle adjustment (BA) based methods. PGO remains a widely adopted back-end in LiDAR SLAM systems [34, 35, 36]. A single-robot pose graph is constructed to refine initial pose estimates from odometry by utilizing the constraints of loop closure. However, traditional PGO frameworks focus on pose consistency rather than directly optimizing point cloud structure, leading to map with poor geometric quality, which limits their effectiveness in generating globally consistent maps for precision-demanding applications.
In contrast, BA-based methods jointly optimize scan poses and scene features by minimizing the geometric residuals of matched primitives, offering improved map accuracy. BALM [23] introduces a closed-form solution for feature parameters to reduce computational complexity, while BALM2 [24] extends this with point clustering and a more efficient second-order solver. HBA [25] adopts a hierarchical BA strategy followed by top-down PGO, significantly improving the quality and consistency of the overall map. PSS-BA [26] introduces quadratic surface modeling for point cloud maps and uses progressive smoothing iteration to optimize map quality. RSO-BA [37] improves robustness via a second-order estimator integrated with a robust kernel function. However, traditional bundle adjustment approaches rely on time-based sliding windows and fail to incorporate long-range spatial constraints introduced by loop closures. These constraints of loop closures, though temporally distant, are spatially adjacent with sharing of the same geometric features and crucial for improving local consistency. To address this gap, we propose a novel BA formulation based on local spatial windows that better leverages spatial correlations for more accurate and consistent mapping.
@@ II-B Multi-Session Map Merging
Outlier rejection constitutes another critical concern in multi-session map merging system. This is because large-scale environments often contain repetitive structures, increasing the risk of incorrect data association and false positive loop closures. Thus, the robust rejection of outliers is essential. RANSAC [40] remains a standard solution by fitting models despite outliers, but its effectiveness diminishes under high outlier ratios and in cases lacking strong priors. PCM [31], adopted in systems such as DCL-SLAM [27] and Disco-SLAM [28], leverages pairwise geometric consistency for loop validation. Although PCM is more robust under uncertain initial estimates, its accuracy can be compromised by unreliable odometry, resulting in over-rejection of valid loop closures.
III Methodology
III-A System Overview
Problem Formulation.
@@ III-A System Overview
In a multi-robot system, the set of pose sequences can be denoted as $S_{N}=\{s_{1},s_{2},\dots,s_{N}\}$, where each sequence $s_{k}$ corresponds to a separate robot. Different sequences (e.g. $s_{i}$ and $s_{j}$), which have different starting points and initial orientations, lack prior information about relative pose transformation. To achieve accurate and effective map fusion, it is essential to estimate the relative transformations between these trajectories and optimize the poses using constraints in overlapping regions to improve geometric consistency.
@@
The problem can therefore be formulated as finding a set of optimized pose sequences $S_{N}^{*}=\{s_{1}^{*},s_{2}^{*},...,s_{N}^{*}\}$, in which all trajectories are aligned to a common coordinate system (typically $s_{1}^{*}$), and the resulting global map exhibits minimal divergence in overlapping regions while maintaining overall consistency.
Framework of Lemon-Mapping.
@@
An overview of the proposed framework is illustrated in Fig. 1, which consists of two main components: the Loop Processing Module (Fig. 1 (a)) detailed in Section III-B and the Map Merging Module (Fig. 1 (b)) described in Section III-C and III-D.
@@
The Loop Processing Module takes the odometry and submaps from multiple agents as input, along with raw loop closure data (including both self-loops and inter-loops) and removes erroneous loops to output the correct loops. The Map Merging Module then uses the correct loops to optimize the multi-robot trajectories through three steps: spatial bundle adjustment and two pose graph optimization (PGO) steps. First, a preliminary global pose estimate is obtained using a conventional loop-based multi-robot pose graph optimization (the first PGO step), which serves as an initial guess and prior information for loop recall. Next, a spatial bundle adjustment is performed within local regions around the loops to fully exploit the geometric constraints from overlapping observations. Finally, the last PGO step refines the trajectories using bundle adjustment constraints and odometry constraints. As the two PGO steps share a similar formulation, they are jointly described in Section III-D, while the spatial bundle adjustment is discussed separately in Section III-C.
III-B Loop Processing Module
@@
The loop processing module is a fundamental component that supports the spatial bundle adjustment described in Section III-C and the final map merging process in Section III-D. It consists of three submodules: Outlier Rejection, Loop Classification, and Loop Recall. This module processes both self-loops and inter-loops, which are initially detected by the robust and SOTA loop detection method RING++ [21].
@@
Figure 2: An example of loop classification, the trajectories of multi-robot and the two types of loops are shown in the figure.
Outlier Rejection.
@@
Loop closure data, including self-loops and inter-loops, may contain incorrect entries, which significantly influences the accuracy of coordinate alignment and pose graph optimization. To solve this problem, we design an outlier rejection method using Generalized Iterative Closest Point (GICP) [41] combined with RANSAC [40].
@@
The process begins with Statistical Outlier Removal (SOR) [42], which filters the noise from the raw point clouds by analyzing the statistical distribution of point-to-neighbor distances. GICP is then employed to align the filtered point clouds, initialized using the transformation estimated by RING++. To validate the alignment, correspondences between the two point clouds are refined using a RANSAC-based correspondence rejection method, which discards outliers inconsistent with rigid transformations. The loop is accepted only if the number of inlier correspondences exceeds a predefined threshold. Additionally, the GICP fitness score is used as a quantitative metric for alignment quality—alignments with poor scores or non-convergence are discarded. This two-stage filtering approach significantly enhances the reliability of loop closures by ensuring that only geometrically valid alignments are retained.
Loop Classification.

### v2

@@ Abstract
Multi-robot collaboration is becoming increasingly critical and presents significant challenges in modern robotics, especially for building a globally consistent, accurate map. Traditional multi-robot pose graph optimization (PGO) methods ensure basic global consistency but ignore the geometric structure of the map, and only use loop closures as constraints between pose nodes, leading to divergence and blurring in overlapping regions. To address this issue, we propose LEMON-Mapping, a loop-enhanced framework for large-scale, multi-session point cloud fusion and optimization. We re-examine the role of loops for multi-robot mapping and introduce three key innovations. First, we develop a robust loop processing mechanism that rejects outliers and a loop recall strategy to recover mistakenly removed but valid loops. Second, we introduce spatial bundle adjustment for multi-robot maps, reducing divergence and eliminating blurring in overlaps. Third, we design a PGO-based approach that leverages refined bundle adjustment constraints to propagate local accuracy to the entire map. We validate LEMON-Mapping on several public datasets and a self-collected dataset. The experimental results show superior mapping accuracy and global consistency of our framework compared to traditional merging methods. Scalability experiments also demonstrate its strong capability to handle scenarios involving numerous robots.
@@ I Introduction
LARGE-SCALE 3D mapping is a fundamental capability in modern robotics, providing rich geometric information that supports tasks such as drone-based inspection [2, 3, 4], autonomous driving [5], and long-term exploration with ground robots [6]. Furthermore, large-scale 3D maps are critical for emerging fields such as embodied AI, where agents interact with complex environments based on spatial understanding [7]. They also play a key role in end-to-end visuomotor navigation systems [8], which rely on accurate environmental priors to enhance generalization and robustness. Especially, multi-robot 3D mapping is essential for large-scale and complex tasks, where robot teams provide greater coverage and robustness compared to a single agent. In scenarios such as search and rescue [9], forest monitoring [10], and subterranean exploration [11], accurate point cloud fusion is vital to realize cooperation among multiple robots under complex and GPS-denied conditions.
@@
Rapid development of LiDAR-based Simultaneous Localization and Mapping (SLAM) techniques [12, 13, 14, 15, 16, 17] has enabled individual robots to perform real-time localization and point cloud map construction. However, most of the existing lightweight SLAM systems [14, 15, 16] for onboard computers are based on direct LiDAR-Inertial-Odometry methods. These methods prioritize pose estimation over map construction to achieve real-time performance. However, the lack of feature extraction and matching in these approaches results in maps with blurred details and poor geometric accuracy. Therefore, the quality of the point cloud maps generated by onboard computers based on lightweight methods requires further improvement.
@@
To address mapping challenges and enhance the geometric accuracy of point cloud maps, various strategies have been proposed. Among them, one category of approaches leverages pose graph optimization (PGO) [18] to handle loop scenarios. Loop closure constraints detected by loop detection methods [19, 20, 21, 22] are added to the pose graph along with odometry constraints to eliminate accumulated drift and improve global consistency of the map. However, PGO-based methods only use the point cloud map to extract descriptors during loop detection and ignore the structure and geometric information. Therefore, although PGO-based methods can improve global consistency based on loop closures, it cannot guarantee an accurate map without divergence and blurring. (We provide examples in Fig. 8 (b) and (c).) In contrast, another type of map refinement method based on bundle adjustment (BA) has demonstrated impressive capabilities to improve the accuracy of point cloud maps in recent years [23, 24, 25, 26]. These approaches utilize the structural characteristics of the point cloud map and improve its quality by minimizing the geometric residuals between feature points and feature structure. Although BA-based methods are capable of generating high-quality maps, their reliance on temporal sequences may prevent them from fully leveraging loop closure information, potentially resulting in poor map consistency. (Examples are illustrated in Fig. 7 (b) and (c).) The dependence on the temporal ordering of poses also limits the applicability of BA-based methods to multi-robot systems.
@@
In addition to the above mapping approaches, to achieve larger-scale map reconstruction and multi-robot collaboration, several multi-robot SLAM systems [27, 28] and offline map fusion techniques [29, 30, 31, 32, 33] have shown promising results in recent years. In general, these methods focus on two aspects: removing outliers in loop closure data and achieving map fusion through multi-robot PGO. However, PGO-based methods are fundamentally the same in single-robot and multi-robot applications as they both use odometry constraints and loop closure constraints for map refinement. As discussed in the previous paragraph, these approaches neither directly utilize the geometric information in multi-robot maps nor fully leverage loop closure data. Consequently, they only ensure the basic alignment and consistency of the global map but fail to generate high-quality point cloud maps. In particular, they often suffer from misalignment in regions with overlap between robots, as demonstrated in Fig. 8 (b) and (c).
@@
In general, PGO-based and BA-based methods cannot guarantee both map accuracy and global consistency in multi-robot scenarios with a large number of loops. Building upon the strengths of existing methods and aiming to address their limitations, we perform an in-depth analysis of previous multi-robot SLAM systems and map merging techniques. The key issue of multi-robot mapping is to solve the divergence and blurring problems of submaps, while maintaining the global consistency of the whole map. We argue that the merging of point cloud maps is essentially a map-driven local point cloud registration problem rather than a pose-dominated pose optimization problem. Building upon this perspective, we identify two essential challenges for achieving effective multi-robot point cloud map fusion: (1) Map fusion should be treated as a map-driven registration problem rather than a pose-centric optimization task such as PGO. (2) Loop closure information should be fully exploited for local registration, not only to ensure global consistency but also to improve local accuracy.
@@
To overcome the above challenge, we propose LEMON-Mapping, a Loop-Enhanced large-scale Multi-session map merging and OptimizatioN framework that achieves globally consistent and geometrically accurate point cloud mapping. Our framework re-examines the ability of loop closure and reasonably enhances its utilization by two aspects. First, we innovatively propose a loop recall mechanism, providing more sufficient information and optimization starting points for subsequent spatial BA.(We provide an example in Fig. 12.) Second, by leveraging sufficient valid loop closures, our proposed spatial BA can effectively deal with multi-robot maps, whereas traditional BA methods fail to address. The spatial BA is performed within a local spatial window around the loop, effectively leveraging multiple observations and abundant constraints from different robots. It directly optimizes the poses from different robots equally within the spatial window; thus, the poses of multiple robots matching the same geometric features (planes, lines) are adjusted simultaneously, generating a locally accurate and consistent map. Although spatial BA improves local map accuracy at loop closures, it lacks the capability for global map fusion. To address this problem, we propose a reasonable PGO-based method which effectively combines local BA constraints and odometry constraints to transfer the local alignment achieved by our spatial BA to the whole map, thereby achieving global consistency and accuracy. We conduct a series of extensive experiments, and the results demonstrate the high mapping accuracy and strong scalability of our multi-session map fusion approach, as shown in Fig. LABEL:fig:headimg and Fig. 14.
@@
  • A scalable multi-session point cloud map merging and optimization framework is designed, which integrates two-step PGO with spatial BA. The method achieves high-precision 3D mapping in large-scale and multi-robot scenarios.
  • A novel spatial BA that can be used for multi-robot mapping is introduced, which operates on loop-based spatial windows to fully utilize loop constraints and jointly optimize multi-robot poses equally. This improves the local accuracy and eliminates the serious map divergence.
  • A pose graph optimization scheme with reasonable sparsification of BA constraints is developed which effectively propagates local accuracy to the global scale, improving both consistency and efficiency.
@@ II Related Works
Figure 1: The framework of our method. Our framework takes multi-robot submaps ,odometry and loop closures as input, and generates accurate and globally consistent merged map. (a) The Loop Processing Module, which removes outliers and recall false negative loops, providing correct and sufficient loops for the spatial BA. (b) The Map Merging Module, which achieves accurate and globally consistent multi-robot map merging through two-step PGO and spatial BA. Each of the three steps in map merging module interacts closely with the multi-robot pose graph.
@@ II-A Single Map Maintenance and Optimization
Single map maintenance and optimization have been extensively studied, with existing approaches falling broadly into two categories: pose graph optimization (PGO) and bundle adjustment (BA) based methods. PGO remains a widely adopted back-end in LiDAR SLAM systems [34, 35, 36]. A single-robot pose graph is constructed to refine initial pose estimates from odometry by utilizing the constraints of loop closures. However, traditional PGO frameworks focus on pose consistency rather than directly optimizing point cloud structure, leading to map with poor geometric quality, which limits their effectiveness in generating globally consistent maps for precision-demanding applications.
In contrast, BA-based methods jointly optimize scan poses and scene features by minimizing the geometric residuals of matched primitives, offering improved map accuracy. BALM [23] introduces a closed-form solution for feature parameters to reduce computational complexity, while BALM2 [24] extends this with point clustering and a more efficient second-order solver. HBA [25] adopts a hierarchical BA strategy followed by top-down PGO, significantly improving the quality and consistency of the overall map. PSS-BA [26] introduces quadratic surface modeling for point cloud maps and uses progressive smoothing iteration to optimize map quality. RSO-BA [37] improves robustness via a second-order estimator integrated with a robust kernel function. However, traditional bundle adjustment approaches rely on time-based sliding windows and fail to incorporate long-range spatial constraints introduced by loop closures. These constraints of loop closures, though temporally distant, are spatially adjacent with sharing of the same geometric features and crucial for improving local consistency. Generally, the existing BA methods mentioned above do not solve this problem well.
@@ II-B Multi-Session Map Merging
Outlier rejection constitutes another critical concern in multi-session map merging system. This is because large-scale environments often contain repetitive structures, increasing the risk of incorrect data association and false positive loop closures. Thus, a robust rejection of outliers is essential. RANSAC [40] remains a standard solution for fitting models despite outliers, but its effectiveness diminishes under high outlier ratios and in cases lacking strong priors. PCM [31], adopted in systems such as DCL-SLAM [27] and Disco-SLAM [28], leverages pairwise geometric consistency for loop validation. Although PCM is more robust under uncertain initial estimates, its accuracy can be compromised by unreliable odometry, resulting in over-rejection of valid loop closures.
III System Overview
III-A Problem Formulation
@@
In a multi-robot system, the set of pose sequences can be denoted as $S_{N}=\{s_{1},s_{2},\dots,s_{N}\}$, where each sequence $s_{k}$ corresponds to a separate robot. Different sequences such as $s_{i}$ and $s_{j}$, which have different starting points and initial orientations, lack prior information on relative pose transformation. To achieve accurate and effective map fusion, it is essential to estimate the relative transformations between these trajectories and optimize the poses using constraints in overlapping regions to improve geometric consistency.
@@
The problem can therefore be formulated as finding a set of optimized pose sequences $S_{N}^{*}=\{s_{1}^{*},s_{2}^{*},...,s_{N}^{*}\}$, such that all trajectories are aligned to a common coordinate system (typically $s_{1}^{*}$), and the resulting global map exhibits minimal divergence in overlapping regions while maintaining overall consistency.
III-B Framework of LEMON-Mapping
@@
An overview of the proposed framework is illustrated in Fig. 1, which is designed for multi-robot point cloud map merging and optimization. Our system consists of two main components: the Loop Processing Module (Fig. 1 (a)) detailed in Section IV and the Map Merging Module (Fig. 1 (b)) described in Section V and Section VI.
@@
The Loop Processing Module takes the odometry and submaps from multiple agents along with raw loop closure data (including both self-loops and inter-loops) as input. It removes erroneous loops to output the correct loops. The Map Merging Module then uses the correct loops to optimize the multi-robot trajectories through three steps: spatial bundle adjustment (BA) and two pose graph optimization (PGO) steps. First, a preliminary global pose estimate is obtained using a conventional multi-robot PGO (FPGO), which serves as an initial guess and prior information for loop recall in the Loop Processing Module. Next, a spatial BA is performed within local regions around the loops to fully exploit the geometric constraints from overlapping observations. Finally, the last PGO (LPGO) step refines the trajectories using constraints of our spatial BA and constraints of odometry. As the two PGO steps share a similar formulation, they are jointly described in Section VI, while the proposed spatial BA is discussed separately in Section V.
IV Loop Processing Module
@@
The Loop Processing Module is a fundamental component that supports the spatial bundle adjustment described in Section V and the two-step pose graph optimization in Section VI. It consists of three submodules: Outlier Rejection, Loop Classification, and Loop Recall. This module processes both self-loops and inter-loops from multiple robots, which are initially detected by the robust and state-of-the-art (SOTA) loop detection method RING++ [21].
@@
Figure 2: An example of loop classification in S3E Campus_1 dataset [41], the trajectories of multi-robot and the two types of loops are shown in the figure.
IV-A Outlier Rejection
@@
Loop closure data may contain incorrect constraints, which seriously affects the accuracy of coordinate alignment and pose graph optimization. To solve this problem, we design an outlier rejection method using the Generalized Iterative Closest Point (GICP) [42] combined with Random Sample Consensus (RANSAC) [40].
@@
The process begins with Statistical Outlier Removal (SOR) [43], which filters the noise from the raw point clouds by analyzing the statistical distribution of point-to-neighbor distances. GICP is then employed to align the filtered point clouds, initialized using the transformation estimated by RING++. To validate the alignment, correspondences between the two point clouds are refined using a RANSAC-based correspondence rejection method, which discards outliers inconsistent with rigid transformations. The loop is accepted only if the number of inlier correspondences exceeds a predefined threshold. Additionally, the GICP fitness score is used as a quantitative metric for alignment quality, and alignments with poor scores or non-convergence are discarded. This two-stage filtering approach significantly enhances the reliability of loop closures by ensuring that only geometrically valid alignments are retained.
IV-B Loop Classification